OpenStudy (anonymous):

how do I integrate sqrt(1+4tan^(2)(theta)) sec^(2)(theta) d(theta)

OpenStudy (anonymous):

Make a substitution to remove the secant term: \(u=\tan \theta\), then \(du=\sec^2\theta\,d\theta\). \[\int\sqrt{1+4\tan^2\theta}\sec^2\theta\,d\theta=\int\sqrt{1+4u^2}\,du\] Make another substitution, introducing a trigonometric form: \(u=\dfrac{1}{2}\tan t\), then \(du=\dfrac{1}{2}\sec^2t\,dt\). \[\int\sqrt{1+4u^2}\,du=\frac{1}{2}\int\sqrt{1+4\left(\frac{1}{2}\tan t\right)^2}\sec^2t\,dt=\frac{1}{2}\int\sqrt{1+\tan^2t}\sec^2t\,dt\] Recall the identity, \(1+\tan^2t=\sec^2t\). Then. \[\frac{1}{2}\int\sqrt{1+\tan^2t}\sec^2t\,dt=\frac{1}{2}\int\sec^3t\,dt\] You can deal with this using integration by parts, or refer to the power reduction formula for secant.

OpenStudy (anonymous):

Of course, you could have made the trig sub right away, setting \(\dfrac{1}{2}\tan t=\tan\theta\), but sometimes subbing one step at a time makes the next step more apparent.